OpenStudy (anonymous):

What is the vertex of the graph of y = –|2x + 4|+ 3? (–2, –3) (–2, 3) (2, –3) (2, 3)

OpenStudy (anonymous):

Solve the two equations: y = -(2x + 4) + 3 2x + y = -1 ----------------1 y = 2x + 4 + 3 2x - y = -7 ---------------------2 Solve these: 2x + y = -1 2x - y = -7 Add them and tell me what do you get ??

OpenStudy (anonymous):

4x-8?

OpenStudy (anonymous):

4x = -8 now divide it by 4 both the sides..

OpenStudy (anonymous):

x=-2

OpenStudy (anonymous):

so itll be -2,3?

OpenStudy (anonymous):

Yes absolutely right..